# Computing the gradients with respect to all parameters of a quantum neural network using a single circuit

### v3

@@ Abstract
Finding gradients is a crucial step in training machine learning models. For quantum neural networks, computing gradients using the parameter-shift rule requires evaluating the cost function twice for each adjustable parameter in the network. When the total number of parameters is large, the quantum circuit must be repeatedly adjusted and executed, leading to significant computational overhead. In this work, we propose a novel approach that computes all gradients using only a single circuit, significantly reducing both the circuit depth and the number of classical registers required. We experimentally validate our approach on both quantum simulators and IBM’s real quantum hardware, demonstrating that our method significantly reduces circuit compilation time compared to the conventional approach, resulting in a substantial speedup in total runtime.
@@ 1 Introduction
In this paper, we will propose an approach in which the gradients with respect to all parameters of a quantum neural network can be computed simultaneously using a single circuit only. The main purpose is to break the limit on the scale of computation on real quantum hardware. As it is well-known, the number of adjustable parameters of a powerful neural network is generally very high. If we have to run two quantum circuits to compute the gradient with respect to a single parameter, then the total number of unique executed circuits could easily exceed the capacity of any real quantum computer in existence. Of course, the execution of these circuits can be broken down into many “jobs”. But it is obviously inconvenient when we are using quantum cloud computation platforms. Especially, if the jobs are submitted separately, then there could be a long waiting time between the execution of the jobs. For example, for free users on the IBM Quantum Experience online platform, this waiting time could vary from $10$ minutes to $4$ hours. For paying users, as also pointed out in Ref. [7], “quantum hardware devices are typically billed and queued per unique circuit”. According to Amazon Braket, the quantum computation platforms Rigetti, IonQ, OQC, and Quera all charge \$0.3 USD per unique circuit in addition to other fees [8]. A conventional method for reducing the number of unique circuits is to “stack” many circuits into a single job. In the Qiskit software development kit for running IBM quantum computers and simulators with Python programs, this can be done using the “append” function [9]. But there is also a restriction on its usage on real quantum computers. For example, on the “ibmq_quito” backend, the number of circuits allowed to be appended is limited to $100$. To break this limit, the Qiskit’s “compose” function can serve as an alternative. According to our experience, it runs much slower than the “append” method does. On platforms that charge by the runtime, slower programs mean higher cost. As of now, the rate in IBM Cloud’s Pay-As-You-Go Plan is \$1.60 USD per second [10]. Nevertheless, it can indeed stack more circuits into one single job so that the gradients with respect to more parameters can be computed together.
@@
In the following, we will use the direct stacking method based on the “compose” function (referred to as the “conventional approach”) as a baseline to compare with the performance of our proposed method that uses a single circuit only (referred to as the “improved approach”). In the next section, we will briefly introduce the typical architecture of quantum neural networks. The parameter-shift rule will be reviewed in Section 3. Subsequently, our improved approach will be presented in Section 4, with its theoretical advantages — namely, reduced circuit depth and lower classical memory requirements — elaborated in Section 5. However, this approach also carries a theoretical disadvantage: the circuit requires more shots to run compared to the conventional approach, as explained in Section 6. To demonstrate that the advantages outweigh this disadvantage, we conduct experiments, and the results are reported in Section 7. These results show a significant speedup in total runtime, particularly as the number of input data points and the scale of the quantum circuit increase. Finally, in Section 8, we discuss the generalization of our approach and highlight the importance of implementing the IF gate on real quantum hardware.
@@ 4 Our improved approach
For this purpose, we add two additional control qubits $q_{a}$ and $q_{b}$ to the circuit, regardless the number of qubits and parameters in the original ansatz. Then for each existing single-parameter rotation gate $RY(\theta_{i})$ ($i=1,...,n$), we added two blocks acting on the qubit it rotates and the above two additional control qubits, to turn the parameter from $\theta_{i}$ to $\theta_{i}+s$ and $\theta_{i}-s$, respectively. Each block contains $2$ controlled-RY gates, $1$ CX (two-qubit controlled-NOT) gate, $1$ measurement and $1$ reset operation. Fig.3 showcases an example of the resultant circuit. For simplicity, only three RY gates in the RealAmplitudes ansatz of the original VQC (corresponding to the three RY gates before any of the red dashed boxes in Fig.1(b)) are studied, and shown with the two additional blocks added to each of them. We also uploaded the diagram of the complete quantum circuit for a VQC with $10$ qubits (which will be used in the Experiments section for the classification task of the MNIST dataset) to Github (available at: MNIST classification.emf).
@@ 6 Theoretical disadvantages
Also, in our approach whether an additional block will be activated is controlled by the measurement result of the qubit $q_{b}$, where quantum uncertainty takes effect. Consequently, executing the circuit for $s^{\prime}=s(2n+1)$ shots does not mean that each cost function will be calculated for $s$ shots exactly. Some statistical fluctuations are inevitable. Nevertheless, the relationship between the number of shots and the precision of the results is also statistical and subjected to quantum uncertainty. A fluctuation of the precision cannot be avoided even if all individual shot numbers are strictly aligned to the same value. Thus, we do not see the need to take extra efforts trying to level off the individual shot numbers of all the cost functions so that they equal exactly to each other.
@@ 7.4 Results
Finally, the impact of noise in the circuits is studied and shown in Fig.6, where the value of the average deviation $\bar{\Delta}$ is calculated as follows. The exact values $f_{i}^{exa}$ ($i=1,...,2n+1$) of all the shifted and unshifted cost functions were calculated numerically beforehand. Then the values $f_{i}^{sim}$ of these cost functions are computed using the noise model in the simulator, where the error rate $\epsilon$ of all quantum gates are taken to be equal for simplicity. The number of shots for each cost function is $s=1000$ (i.e., $s^{\prime}=1000(2n+1)$ for our improved approach). Define
@@ 8 Discussion
(i.e., the gates covered by the original parameter-shift rule [6]), it can also be modified to compute the derivatives for any other quantum gate that requires the general parameter-shift rules [7]. For example, suppose that the first parameterized gate $RY(\theta_{1})$ on the qubit $q_{1}$ in Fig.3 is replaced by a more general gate $U^{\prime}(\theta_{1})$ that does not satisfy Eq. (19). To compute $U^{\prime}(\theta+s)$, all we need is to replace the controlled-RY gate $CRY(s)$ (controlled by $q_{b}$ and acting on $q_{1}$) in the first block right behind $RY(\theta_{1})$ in Fig.3 by two controlled-gates $(U^{\prime}(\theta_{1}))^{-1}$ and $U^{\prime}(\theta_{1}+s)$ in serial, both of which are also controlled by $q_{b}$ and acting on $q_{1}$. Here $(U^{\prime}(\theta_{1}))^{-1}$ is the reverse operation of $U^{\prime}(\theta_{1})$ such that $(U^{\prime}(\theta_{1}))^{-1}U^{\prime}(\theta_{1})=I$. We can see that once the measurement on $q_{b}$ before them (i.e., the measurement operator in the first block of Fig.3) results in $\left|1\right\rangle_{q_{b}}$, then these two gates will be activated. Combining with the original $U^{\prime}(\theta_{1})$ gate, the complete operation on the qubit $q_{1}$ will be $U^{\prime}(\theta_{1}+s)(U^{\prime}(\theta_{1}))^{-1}U^{\prime}(\theta_{1})=U^{\prime}(\theta_{1}+s)$. Thus, the shift of such a general parameterized quantum gate can also be computed with our approach.

### v4

@@ Abstract
Finding gradients is a crucial step in training machine learning models. For quantum neural networks, computing gradients using the parameter-shift rule requires calculating the cost function twice for each adjustable parameter in the network. When the total number of parameters is large, the quantum circuit must be repeatedly adjusted and executed, leading to significant computational overhead. Here we propose an approach to compute all gradients using a single circuit only, significantly reducing both the circuit depth and the number of classical registers required. We experimentally validate our approach on both quantum simulators and IBM’s real quantum hardware, demonstrating that our method significantly reduces circuit compilation time compared to the conventional approach, resulting in a substantial speedup in total runtime.
@@ 1 Introduction
In this paper, we will propose an approach in which the gradients with respect to all parameters of a quantum neural network can be computed simultaneously using a single circuit only. The main purpose is to break the limit on the scale of computation on real quantum hardware. As it is well-known, the number of adjustable parameters of a powerful neural network is generally very high. If we have to run two quantum circuits to compute the gradient with respect to a single parameter, then the total number of unique executed circuits could easily exceed the capacity of any real quantum computer in existence. Of course, the execution of these circuits can be broken down into many “jobs”. But it is obviously inconvenient when we are using quantum cloud computation platforms. Especially, if the jobs are submitted separately, then there could be a long waiting time between the execution of the jobs. For example, for free users on the IBM Quantum Experience online platform, this waiting time could vary from $10$ minutes to $4$ hours. For paying users, as also pointed out in Ref. [7], “quantum hardware devices are typically billed and queued per unique circuit”. According to Amazon Braket, the quantum computation platforms Rigetti, IonQ, OQC, and Quera all charge \$0.3 USD per unique circuit in addition to other fees [8]. A conventional method for reducing the number of unique circuits is to “stack” many circuits into a single job. In the Qiskit software development kit for running IBM quantum computers and simulators with Python programs, this can be done using the “append” function [9]. But there is also a restriction on its usage on real quantum computers. For example, on the “ibmq_quito” backend, the number of circuits allowed to be appended is limited to $100$. To break this limit, the Qiskit’s “compose” function can serve as an alternative. According to our experience, it runs much slower than the “append” method does. On platforms that charge by the runtime, slower programs mean higher cost. As of now, the rate in IBM Cloud’s Pay-As-You-Go Plan is \$96 USD per minute [10]. Nevertheless, it can indeed stack more circuits into one single job so that the gradients with respect to more parameters can be computed together.
@@
In the following, we will use the direct stacking method based on the “compose” function (referred to as the “conventional approach”) as a baseline to compare with the performance of our proposed method that uses a single circuit only (referred to as the “improved approach”). In the next section, we will briefly introduce the typical structure of quantum neural networks. The parameter-shift rule will be reviewed in Section 3. Subsequently, our improved approach will be presented in Section 4, with its theoretical advantages — namely, reduced circuit depth and lower classical memory requirements — elaborated in Section 5. However, this approach also carries some theoretical disadvantages: the circuit needs two additional qubits, and requires more shots to run compared to the conventional approach, as explained in Section 6. To demonstrate that the advantages outweigh these disadvantages, we conduct experiments, and the results are reported in Section 7. These results show a significant speedup in total runtime, particularly as the number of input data points and the scale of the quantum circuit increase. Finally, in Section 8, we discuss the generalization of our approach, and suggest some improvements on real quantum hardware which can make our approach even faster.
@@ 4 Our improved approach
For this purpose, we add two additional control qubits $q_{a}$ and $q_{b}$ to the circuit, regardless the number of qubits and parameters in the original ansatz. Then for each existing single-parameter rotation gate $RY(\theta_{i})$ ($i=1,...,n$), we added two blocks acting on the qubit it rotates and the above two additional control qubits, to turn the parameter from $\theta_{i}$ to $\theta_{i}+s$ and $\theta_{i}-s$, respectively. Each block contains $2$ controlled-RY gates, $1$ CX (two-qubit controlled-NOT) gate, $1$ measurement and $1$ reset operation. Fig.3 showcases an example of the resultant circuit. For simplicity, only three RY gates in the RealAmplitudes ansatz of the original VQC (corresponding to the three RY gates before any of the red dashed boxes in Fig.1(b)) are studied, and shown with the two additional blocks added to each of them. We also uploaded the diagram of the complete quantum circuit for a VQC with $10$ qubits (which will be used in the Experiments section for the classification task of the MNIST dataset) to Github (available at: https:// github.com/gphehub/grad2210/blob/main/ Fig.7_Complete circuit for MNIST classification.emf).
@@ 6 Theoretical disadvantages
Also, in our approach whether an additional block will be activated is controlled by the measurement result of the qubit $q_{b}$, where quantum uncertainty takes effect. Consequently, executing the circuit for $s^{\prime}=s(2n+1)$ shots does not mean that each cost function will be calculated for $s$ shots exactly. Some statistical fluctuations are inevitable. Nevertheless, the relationship between the number of shots and the precision of the results is also statistical and subjected to quantum uncertainty. A fluctuation of the precision cannot be avoided even if all individual shot numbers are strictly aligned to the same value. Thus, we do not see the need to take extra efforts trying to level off the individual shot numbers of all the cost functions so that they equal exactly to each other. (In fact, this can simply be accomplished by discarding the data of some of the shots. But like we said, it seems unnecessary so that our experiments do not include this treatment.)
@@ 7.4 Results
Finally, the impact of noise in the circuits is studied and shown in Fig.6, where the value of the average deviation $\bar{\Delta}$ is calculated as follows. The exact values $f_{i}^{exa}$ ($i=1,...,2n+1$) of all the shifted and unshifted cost functions were calculated numerically beforehand. Then the values $f_{i}^{sim}$ of these cost functions are computed using the noise model in the simulator, where the error rate $\epsilon$ of all quantum gates are taken to be equal for simplicity, and ranges from 0.001 to 0.01. The number of shots for each cost function is $s=1000$ (i.e., $s^{\prime}=1000(2n+1)$ for our improved approach). Define
@@ 8 Discussion
(i.e., the gates covered by the original parameter-shift rule [6]), it can also be modified to compute the gradients for any other quantum gate that requires the general parameter-shift rules [7]. For example, suppose that the first parameterized gate $RY(\theta_{1})$ on the qubit $q_{1}$ in Fig.3 is replaced by a more general gate $U^{\prime}(\theta_{1})$ that does not satisfy Eq. (19). To compute $U^{\prime}(\theta+s)$, all we need is to replace the controlled-RY gate $CRY(s)$ (controlled by $q_{b}$ and acting on $q_{1}$) in the first block right behind $RY(\theta_{1})$ in Fig.3 by two controlled-gates $(U^{\prime}(\theta_{1}))^{-1}$ and $U^{\prime}(\theta_{1}+s)$ in serial, both of which are also controlled by $q_{b}$ and acting on $q_{1}$. Here $(U^{\prime}(\theta_{1}))^{-1}$ is the reverse operation of $U^{\prime}(\theta_{1})$ such that $(U^{\prime}(\theta_{1}))^{-1}U^{\prime}(\theta_{1})=I$. We can see that once the measurement on $q_{b}$ before them (i.e., the measurement operator in the first block of Fig.3) results in $\left|1\right\rangle_{q_{b}}$, then these two gates will be activated. Combining with the original $U^{\prime}(\theta_{1})$ gate, the complete operation on the qubit $q_{1}$ will be $U^{\prime}(\theta_{1}+s)(U^{\prime}(\theta_{1}))^{-1}U^{\prime}(\theta_{1})=U^{\prime}(\theta_{1}+s)$. Thus, the shift of such a general parameterized quantum gate can also be computed with our approach.